### v2

# Search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$

Search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$
M. Ablikim1, M. N. Achasov4,c, P. Adlarson75, O. Afedulidis3, X. C. Ai80, R. Aliberti35, A. Amoroso74A,74C, M. R. An39, Q. An71,58,a, Y. Bai57, O. Bakina36, I. Balossino29A, Y. Ban46,h, H.-R. Bao63, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi74A,74C, E. Bianco74A,74C, A. Bortone74A,74C, I. Boyko36, R. A. Briere5, A. Brueggemann68, H. Cai76, X. Cai1,58, A. Calcaterra28A, G. F. Cao1,63, N. Cao1,63, S. A. Cetin62A, J. F. Chang1,58, G. R. Che43, G. Chelkov36,b, C. Chen43, C. H. Chen9, Chao Chen55, G. Chen1, H. S. Chen1,63, H. Y. Chen20, M. L. Chen1,58,63, S. J. Chen42, S. L. Chen45, S. M. Chen61, T. Chen1,63, X. R. Chen31,63, X. T. Chen1,63, Y. B. Chen1,58, Y. Q. Chen34, Z. J. Chen25,i, Z. Y. Chen1,63, S. K. Choi10A, G. Cibinetto29A, F. Cossio74C, J. J. Cui50, H. L. Dai1,58, J. P. Dai78, A. Dbeyssi18, R.  E. de Boer3, D. Dedovich36, C. Q. Deng72, Z. Y. Deng1, A. Denig35, I. Denysenko36, M. Destefanis74A,74C, F. De Mori74A,74C, B. Ding66,1, X. X. Ding46,h, Y. Ding34, Y. Ding40, J. Dong1,58, L. Y. Dong1,63, M. Y. Dong1,58,63, X. Dong76, M. C. Du1, S. X. Du80, Z. H. Duan42, P. Egorov36,b, Y. H. Fan45, J. Fang1,58, J. Fang59, S. S. Fang1,63, W. X. Fang1, Y. Fang1, Y. Q. Fang1,58, R. Farinelli29A, L. Fava74B,74C, F. Feldbauer3, G. Felici28A, C. Q. Feng71,58, J. H. Feng59, Y. T. Feng71,58, M. Fritsch3, C. D. Fu1, J. L. Fu63, Y. W. Fu1,63, H. Gao63, X. B. Gao41, Y. N. Gao46,h, Yang Gao71,58, S. Garbolino74C, I. Garzia29A,29B, L. Ge80, P. T. Ge76, Z. W. Ge42, C. Geng59, E. M. Gersabeck67, A. Gilman69, K. Goetzen13, L. Gong40, W. X. Gong1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco74A,74C, M. H. Gu1,58, Y. T. Gu15, C. Y. Guan1,63, Z. L. Guan22, A. Q. Guo31,63, L. B. Guo41, M. J. Guo50, R. P. Guo49, Y. P. Guo12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han63, T. T. Han1, X. Q. Hao19, F. A. Harris65, K. K. He55, K. L. He1,63, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng1,58,63, C. Herold60, T. Holtmann3, P. C. Hong34, G. Y. Hou1,63, X. T. Hou1,63, Y. R. Hou63, Z. L. Hou1, B. Y. Hu59, H. M. Hu1,63, J. F. Hu56,j, S. L. Hu12,g, T. Hu1,58,63, Y. Hu1, G. S. Huang71,58, K. X. Huang59, L. Q. Huang31,63, X. T. Huang50, Y. P. Huang1, T. Hussain73, F. Hölzken3, N Hüsken27,35, N. in der Wiesche68, J. Jackson27, S. Janchiv32, J. H. Jeong10A, Q. Ji1, Q. P. Ji19, W. Ji1,63, X. B. Ji1,63, X. L. Ji1,58, Y. Y. Ji50, X. Q. Jia50, Z. K. Jia71,58, D. Jiang1,63, H. B. Jiang76, P. C. Jiang46,h, S. S. Jiang39, T. J. Jiang16, X. S. Jiang1,58,63, Y. Jiang63, J. B. Jiao50, J. K. Jiao34, Z. Jiao23, S. Jin42, Y. Jin66, M. Q. Jing1,63, X. M. Jing63, T. Johansson75, S. Kabana33, N. Kalantar-Nayestanaki64, X. L. Kang9, X. S. Kang40, M. Kavatsyuk64, B. C. Ke80, V. Khachatryan27, A. Khoukaz68, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui1,63, N.  Kumar26, A. Kupsc44,75, W. Kühn37, J. J. Lane67, P.  Larin18, L. Lavezzi74A,74C, T. T. Lei71,58, Z. H. Lei71,58, M. Lellmann35, T. Lenz35, C. Li47, C. Li43, C. H. Li39, Cheng Li71,58, D. M. Li80, F. Li1,58, G. Li1, H. B. Li1,63, H. J. Li19, H. N. Li56,j, Hui Li43, J. R. Li61, J. S. Li59, Ke Li1, L. J Li1,63, L. K. Li1, Lei Li48, M. H. Li43, P. R. Li38,l, Q. M. Li1,63, Q. X. Li50, R. Li17,31, S. X. Li12, T.  Li50, W. D. Li1,63, W. G. Li1,a, X. Li1,63, X. H. Li71,58, X. L. Li50, X. Z. Li59, Xiaoyu Li1,63, Y. G. Li46,h, Z. J. Li59, Z. X. Li15, C. Liang42, H. Liang71,58, H. Liang1,63, Y. F. Liang54, Y. T. Liang31,63, G. R. Liao14, L. Z. Liao50, J. Libby26, A.  Limphirat60, C. C. Lin55, D. X. Lin31,63, T. Lin1, B. J. Liu1, B. X. Liu76, C. Liu34, C. X. Liu1, F. H. Liu53, Fang Liu1, Feng Liu6, G. M. Liu56,j, H. Liu38,k,l, H. B. Liu15, H. M. Liu1,63, Huanhuan Liu1, Huihui Liu21, J. B. Liu71,58, J. Y. Liu1,63, K. Liu38,k,l, K. Y. Liu40, Ke Liu22, L. Liu71,58, L. C. Liu43, Lu Liu43, M. H. Liu12,g, P. L. Liu1, Q. Liu63, S. B. Liu71,58, T. Liu12,g, W. K. Liu43, W. M. Liu71,58, X. Liu39, X. Liu38,k,l, Y. Liu80, Y. Liu38,k,l, Y. B. Liu43, Z. A. Liu1,58,63, Z. D. Liu9, Z. Q. Liu50, X. C. Lou1,58,63, F. X. Lu59, H. J. Lu23, J. G. Lu1,58, X. L. Lu1, Y. Lu7, Y. P. Lu1,58, Z. H. Lu1,63, C. L. Luo41, M. X. Luo79, T. Luo12,g, X. L. Luo1,58, X. R. Lyu63, Y. F. Lyu43, F. C. Ma40, H. Ma78, H. L. Ma1, J. L. Ma1,63, L. L. Ma50, M. M. Ma1,63, Q. M. Ma1, R. Q. Ma1,63, T. Ma71,58, X. T. Ma1,63, X. Y. Ma1,58, Y. Ma46,h, Y. M. Ma31, F. E. Maas18, M. Maggiora74A,74C, S. Malde69, Y. J. Mao46,h, Z. P. Mao1, S. Marcello74A,74C, Z. X. Meng66, J. G. Messchendorp13,64, G. Mezzadri29A, H. Miao1,63, T. J. Min42, R. E. Mitchell27, X. H. Mo1,58,63, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie20, I. B. Nikolaev4,c, Z. Ning1,58, S. Nisar11,m, Q. L. Niu38,k,l, W. D. Niu55, Y. Niu 50, S. L. Olsen63, Q. Ouyang1,58,63, S. Pacetti28B,28C, X. Pan55, Y. Pan57, A.  Pathak34, P. Patteri28A, Y. P. Pei71,58, M. Pelizaeus3, H. P. Peng71,58, Y. Y. Peng38,k,l, K. Peters13,e, J. L. Ping41, R. G. Ping1,63, S. Plura35, V. Prasad33, F. Z. Qi1, H. Qi71,58, H. R. Qi61, M. Qi42, T. Y. Qi12,g, S. Qian1,58, W. B. Qian63, C. F. Qiao63, X. K. Qiao80, J. J. Qin72, L. Q. Qin14, L. Y. Qin71,58, X. S. Qin50, Z. H. Qin1,58, J. F. Qiu1, Z. H. Qu72, C. F. Redmer35, K. J. Ren39, A. Rivetti74C, M. Rolo74C, G. Rong1,63, Ch. Rosner18, S. N. Ruan43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning75, M. Scodeggio29A, K. Y. Shan12,g, W. Shan24, X. Y. Shan71,58, Z. J Shang38,k,l, J. F. Shangguan55, L. G. Shao1,63, M. Shao71,58, C. P. Shen12,g, H. F. Shen1,8, W. H. Shen63, X. Y. Shen1,63, B. A. Shi63, H. Shi71,58, H. C. Shi71,58, J. L. Shi12,g, J. Y. Shi1, Q. Q. Shi55, S. Y. Shi72, X. Shi1,58, J. J. Song19, T. Z. Song59, W. M. Song34,1, Y.  J. Song12,g, Y. X. Song46,h,n, S. Sosio74A,74C, S. Spataro74A,74C, F. Stieler35, Y. J. Su63, G. B. Sun76, G. X. Sun1, H. Sun63, H. K. Sun1, J. F. Sun19, K. Sun61, L. Sun76, S. S. Sun1,63, T. Sun51,f, W. Y. Sun34, Y. Sun9, Y. J. Sun71,58, Y. Z. Sun1, Z. Q. Sun1,63, Z. T. Sun50, C. J. Tang54, G. Y. Tang1, J. Tang59, M. Tang71,58, Y. A. Tang76, L. Y. Tao72, Q. T. Tao25,i, M. Tat69, J. X. Teng71,58, V. Thoren75, W. H. Tian59, Y. Tian31,63, Z. F. Tian76, I. Uman62B, Y. Wan55, S. J. Wang 50, B. Wang1, B. L. Wang63, Bo Wang71,58, D. Y. Wang46,h, F. Wang72, H. J. Wang38,k,l, J. J. Wang76, J. P. Wang 50, K. Wang1,58, L. L. Wang1, M. Wang50, Meng Wang1,63, N. Y. Wang63, S. Wang38,k,l, S. Wang12,g, T.  Wang12,g, T. J. Wang43, W. Wang59, W.  Wang72, W. P. Wang35,71,o, X. Wang46,h, X. F. Wang38,k,l, X. J. Wang39, X. L. Wang12,g, X. N. Wang1, Y. Wang61, Y. D. Wang45, Y. F. Wang1,58,63, Y. L. Wang19, Y. N. Wang45, Y. Q. Wang1, Yaqian Wang17, Yi Wang61, Z. Wang1,58, Z. L.  Wang72, Z. Y. Wang1,63, Ziyi Wang63, D. H. Wei14, F. Weidner68, S. P. Wen1, Y. R. Wen39, U. Wiedner3, G. Wilkinson69, M. Wolke75, L. Wollenberg3, C. Wu39, J. F. Wu1,8, L. H. Wu1, L. J. Wu1,63, X. Wu12,g, X. H. Wu34, Y. Wu71,58, Y. H. Wu55, Y. J. Wu31, Z. Wu1,58, L. Xia71,58, X. M. Xian39, B. H. Xiang1,63, T. Xiang46,h, D. Xiao38,k,l, G. Y. Xiao42, S. Y. Xiao1, Y.  L. Xiao12,g, Z. J. Xiao41, C. Xie42, X. H. Xie46,h, Y. Xie50, Y. G. Xie1,58, Y. H. Xie6, Z. P. Xie71,58, T. Y. Xing1,63, C. F. Xu1,63, C. J. Xu59, G. F. Xu1, H. Y. Xu66, M. Xu71,58, Q. J. Xu16, Q. N. Xu30, W. Xu1, W. L. Xu66, X. P. Xu55, Y. C. Xu77, Z. P. Xu42, Z. S. Xu63, F. Yan12,g, L. Yan12,g, W. B. Yan71,58, W. C. Yan80, X. Q. Yan1, H. J. Yang51,f, H. L. Yang34, H. X. Yang1, Tao Yang1, Y. Yang12,g, Y. F. Yang43, Y. X. Yang1,63, Yifan Yang1,63, Z. W. Yang38,k,l, Z. P. Yao50, M. Ye1,58, M. H. Ye8, J. H. Yin1, Z. Y. You59, B. X. Yu1,58,63, C. X. Yu43, G. Yu1,63, J. S. Yu25,i, T. Yu72, X. D. Yu46,h, Y. C. Yu80, C. Z. Yuan1,63, J. Yuan34, L. Yuan2, S. C. Yuan1, Y. Yuan1,63, Y. J. Yuan45, Z. Y. Yuan59, C. X. Yue39, A. A. Zafar73, F. R. Zeng50, S. H.  Zeng72, X. Zeng12,g, Y. Zeng25,i, Y. J. Zeng59, X. Y. Zhai34, Y. C. Zhai50, Y. H. Zhan59, A. Q. Zhang1,63, B. L. Zhang1,63, B. X. Zhang1, D. H. Zhang43, G. Y. Zhang19, H. Zhang71,58, H. Zhang80, H. C. Zhang1,58,63, H. H. Zhang59, H. H. Zhang34, H. Q. Zhang1,58,63, H. R. Zhang71,58, H. Y. Zhang1,58, J. Zhang59, J. Zhang80, J. J. Zhang52, J. L. Zhang20, J. Q. Zhang41, J. S. Zhang12,g, J. W. Zhang1,58,63, J. X. Zhang38,k,l, J. Y. Zhang1, J. Z. Zhang1,63, Jianyu Zhang63, L. M. Zhang61, Lei Zhang42, P. Zhang1,63, Q. Y. Zhang34, R. Y Zhang38,k,l, Shuihan Zhang1,63, Shulei Zhang25,i, X. D. Zhang45, X. M. Zhang1, X. Y. Zhang50, Y.  Zhang72, Y.  T. Zhang80, Y. H. Zhang1,58, Y. M. Zhang39, Yan Zhang71,58, Yao Zhang1, Z. D. Zhang1, Z. H. Zhang1, Z. L. Zhang34, Z. Y. Zhang43, Z. Y. Zhang76, Z. Z.  Zhang45, G. Zhao1, J. Y. Zhao1,63, J. Z. Zhao1,58, Lei Zhao71,58, Ling Zhao1, M. G. Zhao43, N. Zhao78, R. P. Zhao63, S. J. Zhao80, Y. B. Zhao1,58, Y. X. Zhao31,63, Z. G. Zhao71,58, A. Zhemchugov36,b, B. Zheng72, B. M. Zheng34, J. P. Zheng1,58, W. J. Zheng1,63, Y. H. Zheng63, B. Zhong41, X. Zhong59, H.  Zhou50, J. Y. Zhou34, L. P. Zhou1,63, S.  Zhou6, X. Zhou76, X. K. Zhou6, X. R. Zhou71,58, X. Y. Zhou39, Y. Z. Zhou12,g, J. Zhu43, K. Zhu1, K. J. Zhu1,58,63, K. S. Zhu12,g, L. Zhu34, L. X. Zhu63, S. H. Zhu70, S. Q. Zhu42, T. J. Zhu12,g, W. D. Zhu41, Y. C. Zhu71,58, Z. A. Zhu1,63, J. H. Zou1, J. Zu71,58 (BESIII Collaboration) 1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
64 University of Groningen, NL-9747 AA Groningen, The Netherlands
65 University of Hawaii, Honolulu, Hawaii 96822, USA
66 University of Jinan, Jinan 250022, People’s Republic of China
67 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
68 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
69 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
70 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
71 University of Science and Technology of China, Hefei 230026, People’s Republic of China
72 University of South China, Hengyang 421001, People’s Republic of China
73 University of the Punjab, Lahore-54590, Pakistan
74 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
75 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
76 Wuhan University, Wuhan 430072, People’s Republic of China
77 Yantai University, Yantai 264005, People’s Republic of China
78 Yunnan University, Kunming 650500, People’s Republic of China
79 Zhejiang University, Hangzhou 310027, People’s Republic of China
80 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
Abstract
Using 9.0 $\rm fb^{-1}$ of $e^{+}e^{-}$ collision data collected at center-of-mass energies from 4.178 to 4.278 GeV with the BESIII detector at the BEPCII collider, we perform the first search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$. No signal is observed and the upper limit on the ratio of branching fractions $\mathcal{B}(\chi_{c1}(3872)\to\gamma\psi_{2}(3823),\psi_{2}(3823)\to\gamma\chi_{c1})/\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)$ is set at 0.075 at the 90% confidence level. Our result contradicts theoretical predictions under the assumption that the $\chi_{c1}(3872)$ is the pure charmonium state $\chi_{c1}(2P)$.
I Introduction
As the prototypical example of charmonium-like $XYZ$ states, the $\chi_{c1}(3872)$ has been extensively investigated in the past two decades since it was discovered by the Belle experiment [1] in 2003. From a global fit to the measurements by LHCb, BESIII, Belle, BaBar, and others, its mass and width are determined to be $\rm M=3871.65\pm 0.06$ MeV/$c^{2}$ and $\Gamma=1.19\pm 0.21$ MeV, respectively [2]. Its spin, parity and charge-conjugation parity quantum numbers are determined to be $J^{PC}=1^{++}$ [3]. So far, the observed decay modes of the particle include $D^{*0}\bar{D^{0}}+c.c.$, $\pi^{+}\pi^{-}J/\psi$, $\omega J/\psi$, $\gamma J/\psi$, and $\pi^{0}\chi_{c1}$ [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Although tremendous effort has been made from both the experimental and theoretical sides, the interpretation of the $\chi_{c1}(3872)$ remains inconclusive. Due to the proximity of its mass to the $D^{*0}\bar{D^{0}}+c.c.$ mass threshold, it is conjectured to have a large $D^{*0}\bar{D^{0}}+c.c.$ molecular component [14, 15]. Indeed, some theoretical models consider it to be a mixture of a conventional $2^{3}P_{1}$ charmonium state $\chi_{c1}(2P)$ and a $D^{*0}\bar{D^{0}}+c.c.$ molecule [16, 17].
Measurements of new $\chi_{c1}(3872)$ decay modes can help to improve our understanding of its internal structure. Ref. [18] extracted the absolute branching fractions of the known $\chi_{c1}(3872)$ decays by performing a global fit of the absolute branching fraction of the $B^{+}\to\chi_{c1}(3872)K^{+}$ channel measured by BaBar [19] together with information from other experiments. The fraction of $\chi_{c1}(3872)$ decays not observed in experiments is estimated to be $31.9^{+18.1}_{-31.5}$%. The work is carried out by assuming the $\chi_{c1}(3872)$ has universal properties in different production and decay mechanisms. Meanwhile, Ref. [20] also reported the branching fractions with consideration of the threshold effect of $D^{*0}\bar{D^{0}}+c.c.$ and a possible bound state below the threshold or a virtual state in $B^{+}\to\chi_{c1}(3872)K^{+}$ decay. If the $\chi_{c1}(3872)$ contains a component of the excited spin-triplet state $\chi_{c1}(2P)$, then the radiative decay $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$ could happen naturally via a E1 transition [21], where the $\psi_{2}(3823)$ is considered as the $1^{3}D_{2}$ charmonium state. The BESIII experiment has reported the observation of $e^{+}e^{-}\to\gamma\chi_{c1}(3872)$ at center-of-mass energies $\sqrt{s}=4.178-4.278$ GeV [22, 23]. Using the $\chi_{c1}(3872)$ signal produced in these data samples, we search for the radiative transition $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$, where the $\psi_{2}(3823)$ is reconstructed with the cascade decay $\psi_{2}(3823)\to\gamma\chi_{c1}$, $\chi_{c1}\to\gamma J/\psi$, $J/\psi\to\ell^{+}\ell^{-}$ ($\ell=e,\mu$). The branching fraction ratio of this decay relative to the well-established $\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi$ decay, $\mathcal{R}_{\chi_{c1}(3872)}\equiv\mathcal{B}(\chi_{c1}(3872)\to\gamma\psi_{2}(3823),\psi_{2}(3823)\to\gamma\chi_{c1})/\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)$, is determined in this work.
[…]
Table 1: The calculated values for $\mathcal{R}_{\chi_{c1}(2P)}$, by including as input values the partial decay widths $\Gamma_{\chi_{c1}(2P)\to\gamma\psi(1^{3}D_{2})}$ and $\Gamma_{\psi(1^{3}D_{2})\to\gamma\chi_{c1}(1P)}$ predicted by the NR and GI models and LQCD, the total widths, $\Gamma_{\chi_{c1}(3872)}$ and $\Gamma_{\psi_{2}(3823)}$, and the branching fraction ${\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)}$. The ``"``"``-"` ` - " means unavailable. The two values of the ratio for the LQCD case correspond to the results by taking the $\Gamma_{\chi_{c1}(2P)\to\gamma\psi(1^{3}D_{2})}$ width from the NR and GI models as input, respectively.
$\Gamma_{\chi_{c1}(3872)}=1190\pm 210$ keV [2]
$\Gamma_{\psi_{2}(3823)}=520\pm 100$ keV [24]
$\mathcal{B}(\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi)=(3.8\pm 1.2)\times 10^{-2}$ [2]
NR [21] GI [21] LQCD [24]
$\Gamma_{\chi_{c1}(2P)\to\gamma\psi(1^{3}D_{2})}$ (keV) 35 18 --
$\Gamma_{\psi(1^{3}D_{2})\to\gamma\chi_{c1}(1P)}$ (keV) 307 268 $337\pm 27$
$\mathcal{R}_{\chi_{c1}(2P)}$ $0.46\pm 0.19$ $0.21\pm 0.09$ $0.50\pm 0.21$, $0.26\pm 0.11$
II BESIII Detector and data sets
[…]
The $e^{+}e^{-}$ collision data collected at $\sqrt{s}=4.178-4.278$ GeV is used in this analysis. The integrated luminosity at each energy point is measured with the Bhabha scattering process with a precision better than 1% [29] as listed in Table 2. A geant4-based [30] software package is used to generate the Monte Carlo (MC) simulated data samples. The inclusive MC samples, used to estimate the backgrounds, include the open-charm hadronic processes, continuum processes, and the initial-state-radiation effects, and are generated with kkmc [31] in conjunction with evtgen [32]. The signal MC samples $e^{+}e^{-}\to\gamma\chi_{c1}(3872)$, with the decay chain $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$, $\psi_{2}(3823)\to\gamma\chi_{c1}$, $\chi_{c1}\to\gamma J/\psi$, $J/\psi\to\ell^{+}\ell^{-}$ ($\ell=e,\mu$), are used to determine the detection efficiency. The $e^{+}e^{-}\to\gamma\chi_{c1}(3872)$ decay is simulated as an E1 transition according to the measurement from BESIII [22]. The $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$ and $\psi_{2}(3823)\to\gamma\chi_{c1}$ decays are produced with a phase space model.
Table 2: The data sets and their integrated luminosity at each energy point.
$\sqrt{s}$ (GeV) Luminosity ($\textrm{pb}^{-1}$)
4.178 3189.0
4.189 526.7
4.199 526.0
4.209 517.1
4.219 514.6
4.226 1101.0
4.236 530.3
4.244 538.1
4.258 828.4
4.267 531.1
4.278 175.7
III Event Selection and result
According to the decay chain of the signal process, $e^{+}e^{-}\to\gamma\chi_{c1}(3872)$, $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$, $\psi_{2}(3823)\to\gamma\chi_{c1}$, $\chi_{c1}\to\gamma J/\psi$, $J/\psi\to\ell^{+}\ell^{-}$ ($\ell=e,\mu$), the final state contains a lepton pair from the $J/\psi$ decay and four radiative photons. For the leptons, each corresponding charged track is required to have its point of closest approach to the beam axis within 1 cm in the radial direction and within 10 cm along the beam direction, and to lie within the polar-angle coverage of the MDC, $|\cos\theta|<0.93$, in the laboratory frame. We require exactly two good charged tracks in the candidate events. EMC information discriminates between the electrons and muons: electrons are required to deposit at least 0.8 GeV in the EMC, and the muons less than 0.4 GeV. Photons are reconstructed from isolated showers in the EMC, at least 10 degrees away from any charged track, with an energy deposit of at least 25 MeV in both the barrel ($|\cos\theta|<0.80$) and the end-cap ($0.920.86<|\cos\theta|<$ 0.92) regions. In order to suppress electronic noise unrelated to the event, the EMC time, $t$, of the photon candidate must be in the range $0\leq t\leq 700~{}\rm{ns}$, consistent with collision events. We require at least four photons for each candidate event.
A four-constraint (4C) kinematic fit is applied to constrain the total four-momentum of the lepton pair and the four photons to that of the colliding beams, to suppress backgrounds and improve the resolution. For events with more than four photons, the combination with the best fit quality corresponding to the minimum fit chi-square, $\chi^{2}_{\rm 4C}$, is retained. The $J/\psi$ is reconstructed by requiring the invariant mass, $M(\ell\ell)$, of the lepton pair to satisfy $|M(\ell\ell)-m(J/\psi)|<30$ MeV/$c^{2}$, where $m(J/\psi)$ is the nominal $J/\psi$ mass. The selection criteria are optimized by maximizing the punzi figure-of-merit $S/(\frac{a}{2}+\sqrt{B})$ [33], where the number of signal events ($S$) is determined with the signal MC sample, the background ($B$) is estimated with the inclusive MC, and the expected statistical significance ($a$) is set to be 3. The dominant background is from the process $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$. After the $J/\psi$ selection, we veto $\pi^{0}$ candidates by requiring that the invariant mass of all photon pairs is more than 15 MeV/$c^{2}$ away from the nominal $\pi^{0}$ mass. After these requirements, a seven-constraint (7C) kinematic fit with an additional three constraints on the masses of $M(\ell\ell)$, $M(\gamma\ell\ell)$, and $M(\gamma\gamma\ell\ell)$ to the nominal masses of $J/\psi$, $\chi_{c1}$, and $\psi_{2}(3823)$, respectively, is applied to distinguish the radiative photon in each cascade decay. The best-fit combination with the minimum chi-square, $\chi^{2}_{\rm 7C}$, is retained; $\chi^{2}_{\rm 7C}<100$ is also required to further suppress the combinatorial backgrounds. One possible peaking background is $\psi_{2}(3823)\to\gamma\chi_{c2},~{}\chi_{c2}\to\gamma J/\psi$, the contribution of which is estimated according to the measurement of the branching fraction ratio of $\psi_{2}(3823)\to\gamma\chi_{c2}$ to $\psi_{2}(3823)\to\gamma\chi_{c1}$ in Ref. [34]. The ratio of the yields of $\psi_{2}(3823)\to\gamma\chi_{c2}$ to $\psi_{2}(3823)\to\gamma\chi_{c1,2}$, is about 1.5%, which is taken into account as a source of systematic uncertainty.
Figure 1 shows the distribution of the invariant mass of the radiative photon and the $\psi_{2}(3823)$, $M(\gamma\psi_{2}(3823))$ for the selected candidates, summed over all the energy points. No signal is observed in the $\chi_{c1}(3872)$ signal region in data. The three events around 3.93 GeV are very unlikely to be from the $\chi_{c2}(2P)$ decays since no $\chi_{c2}(2P)$ signal was observed in its more favourable radiative transition to $\psi(2S)$ [9]. After normalizing the MC samples according to the luminosity and cross section in data, the contributions of the $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$ process and of the other backgrounds, estimated with the inclusive MC sample, are also shown in Fig. 1.
Figure 1: Distribution of $M(\gamma\psi_{2}(3823))$. The dots with error bars are data, the red histogram is the signal MC sample with arbitrary scale, the filled blue histogram is the inclusive MC sample without the process $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$, and the green stacked histogram is the contribution from $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$.
The branching ratio $\mathcal{R}_{\chi_{c1}(3872)}$ is calculated as
$\mathcal{R}_{\chi_{c1}(3872)}$ $=\frac{N_{\rm obs}-r\cdot N^{\rm sdb}_{\rm obs}}{N_{\pi^{+}\pi^{-}J/\psi}\cdot\frac{\epsilon_{\gamma\psi_{2}(3823)}}{\epsilon_{\pi^{+}\pi^{-}J/\psi}}\cdot\mathcal{B}(\chi_{c1}\to\gamma J/\psi)},$ (1)
where $N_{\rm obs}=0$ is the number of observed events from all data in the $\chi_{c1}(3872)$ signal region [3.855, 3.885] GeV/$c^{2}$ which covers around ±$\pm 3\sigma$ of the signal shape according to the signal MC distributions, $N_{\rm obs}^{\rm sdb}=4$ is the number of the observed events in the $\chi_{c1}(3872)$ sideband region [3.840, 3.855] and [3.885, 3.940] GeV/$c^{2}$; $r$, the background scaling factor from the sideband regions to the signal region, is 0.474 based on the inclusive MC sample (taking into account its systematic uncertainty; see Sec. IV); $N_{\pi^{+}\pi^{-}J/\psi}=80.7\pm 9.0$ is taken from the BESIII measurement [10]; the branching fraction $\mathcal{B}(\chi_{c1}\to\gamma J/\psi)=0.343\pm 0.010$ is quoted from the PDG [2]; $\epsilon_{\gamma\psi_{2}(3823)}$ is the efficiency for the signal process reconstruction, determined with the signal MC sample; and $\epsilon_{\pi^{+}\pi^{-}J/\psi}$ is the efficiency of the process $\chi_{c1}(3872)\to\pi^{+}\pi^{-}J/\psi$ [10]. The efficiency ratio $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$ at each energy point is shown in Fig. 2, which is almost independent on the center-of-mass energy. The mean value with the standard deviation, $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}=0.433\pm 0.004$, is used to calculate the $\mathcal{R}_{\chi_{c1}(3872)}$ value. The upper limit of $\mathcal{R}_{\chi_{c1}(3872)}$ at the 90% confidence level (C.L.) is computed with the TRolke program implemented in the ROOT framework [35] by assuming the background $N_{\rm obs}^{\rm sdb}$ and the denominator of $\mathcal{R}_{\chi_{c1}(3872)}$ follow Poisson and Gaussian distributions, respectively, where the systematic uncertainties discussed in the following section is taken as the standard deviation of the Gaussian function to be considered in the upper limit. We obtain an upper limit of $\mathcal{R}_{\chi_{c1}(3872)}<0.075$ at the 90% C.L.
Figure 2: Values of $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$ at each energy point (blue dots). The red line indicates the mean value.
IV systematic uncertainties
Systematic uncertainties on $\mathcal{R}_{\chi_{c1}(3872)}$ arise mainly from the estimations of $r$, the possible peaking background of $\psi_{2}(3823)\to\gamma\chi_{c2}\to\gamma\gamma J/\psi$, $N_{\pi^{+}\pi^{-}J/\psi}$, $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$, and $\mathcal{B}(\chi_{c1}\to\gamma J/\psi)$. The background scaling factor $r$ is determined from the inclusive MC samples including the process $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$. We use a 1st or 2nd order polynomial function to fit the $M(\gamma\psi_{2}(3823))$ distribution from the inclusive MC samples; the $r$ value is calculated several times using the parameters from the fit and varying them within $1\sigma$. The value $r=0.474$ is chosen from the obtained values since it provides the most conservative upper limit. The contribution of the potential peaking background of $\psi_{2}(3823)\to\gamma\chi_{c2}\to\gamma\gamma J/\psi$ is estimated with the related measurements mentioned previously within one errors, and the result providing the most conservative $\mathcal{R}_{\chi_{c1}(3872)}$ upper limit is retained.
Both statistical and systematic uncertainties of $N_{\pi^{+}\pi^{-}J/\psi}$ contribute as sources of systematic uncertainty, where the statistical part (11.2%) is obtained by assuming that $N_{\pi^{+}\pi^{-}J/\psi}$ follows a Poisson distribution, and the systematic part (4.1%) is obtained from Ref. [10] where the dominant contribution is from the parametrization of the $\chi_{c1}(3872)$ signal shape. The systematic uncertainty (2.9%) due to $\mathcal{B}(\chi_{c1}\to\gamma J/\psi)$ is taken from the PDG [2]. The systematic uncertainty of $\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$ comes mainly from the tracking (2.0%), the photon selection (3.0%), and the kinematic fit (2.2%) uncertainties, estimated with the control sample $e^{+}e^{-}\to\pi^{0}\pi^{0}J/\psi$. The systematic uncertainty due to the $\pi^{0}$ veto is mainly caused by potential differences in the angular distributions of the radiative photon between the data and the signal MC sample, and it is estimated by changing the angular distribution of the radiative $\gamma$ in $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$ to $1\pm\cos^{2}\theta$ (from flat) in the generator model. The relative difference of 5.3% between the efficiencies obtained with the photon angular distributions of $1-\rm cos^{2}\theta$ and $1+\cos^{2}\theta$ is taken as the systematic uncertainty.
The systematic uncertainties are listed in Table 3. The total systematic uncertainty is obtained by summing all systematic uncertainties in quadrature, assuming they are uncorrelated.
Table 3: The relative systematic uncertainties on $\mathcal{R}_{\chi_{c1}(3872)}$. Systematics on the sideband scaling ratio, $r$, are treated separately (see text).
Item sys. (%)
$N_{\pi^{+}\pi^{-}J/\psi}$ stat. 11.2
sys. 4.1
tracking 2.0
$\epsilon_{\gamma\psi_{2}(3823)}/\epsilon_{\pi^{+}\pi^{-}J/\psi}$ photon 3.0
kinematic fit 2.2
$\pi^{0}$ veto 5.3
$\mathcal{B}(\chi_{c1}\to\gamma J/\psi)$ 2.9
Sum 14.1
V summary
In summary, we search for the radiative decay $\chi_{c1}(3872)\to\gamma\psi_{2}(3823)$ for the first time by using the $e^{+}e^{-}$ collision data accumulated at $\sqrt{s}=4.178-4.278$ GeV with the BESIII detector. No signal is observed, and the upper limit on the branching fraction ratio $\mathcal{R}_{\chi_{c1}(3872)}$ is determined to be 0.075 at the 90% C.L. This upper limit is more than $1\sigma$ below the theoretical calculations of $\mathcal{R}_{\chi_{c1}(3872)}$ under the assumption that the $\chi_{c1}(3872)$ is the pure charmonium state $\chi_{c1}(2P)$, listed in Table 1, and much smaller than the predictions based on the FWHMs measured by LHCb and BESIII [25, 26]. Our result therefore indicates that the $\chi_{c1}(3872)$ is not a pure $\chi_{c1}(2P)$ charmonium state.
Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2020YFA0406300, 2020YFA0406400; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11805090, 11635010, 11735014, 11835012, 11935015, 11935016, 11935018, 11961141012, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; CAS Key Research Program of Frontier Sciences under Contracts Nos. QYZDJ-SSW-SLH003, QYZDJ-SSW-SLH040; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; European Union’s Horizon 2020 research and innovation programme under Marie Sklodowska-Curie grant agreement under Contract No. 894790; German Research Foundation DFG under Contracts Nos. 455635585, Collaborative Research Center CRC 1044, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.